Direct Materials and Direct Labor Variance Analysis

Faucet Industries Inc. manufactures faucets in a small manufacturing facility. The faucets are made from zinc. Faucet Industries has 40 employees. Each employee presently provides 35 hours of labor per week. Information about a production week is as follows:

Standard wage per hour $13.80
Standard labor time per faucet 15 min.
Standard number of lbs. of zinc 1.8 lbs.
Standard price per lb. of zinc $10.50
Actual price per lb. of zinc $10.75
Actual lbs. of zinc used during the week 15,000 lbs.
Number of faucets produced during the week 8,100
Actual wage per hour $14.20
Actual hours per week 1,400 hrs.

Required:

a. Determine the standard cost per faucet for direct materials and direct labor. Do not round intermediate calculations and round your answers to the nearest cent.

Direct materials standard cost per faucet $
Direct labor standard cost per faucet   
Total cost per faucet $

b. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative amount, and an unfavorable variance as a positive amount

Price variance $
Quantity variance   
Total direct materials cost variance $

c. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative amount, and an unfavorable variance as a positive amount.

Rate variance $
Time variance   
Total direct labor cost variance $
Answer #1
a
Direct materials standard cost per faucet 18.90 =1.8*10.5
Direct labor standard cost per faucet 3.45 =13.8*(15/60)
Total cost per faucet 22.35
b
Price variance 3750 Unfavorable =15000*(10.75-10.5)
Quantity variance 4410 Unfavorable =10.5*(15000-8100*1.8)
Total direct materials cost variance 8160 Unfavorable =(15000*10.75)-(8100*1.8*10.5)
c
Rate variance 560 Unfavorable =1400*(14.2-13.8)
Time variance -8625 Favorable =13.8*(1400-8100*15/60)
Total direct labor cost variance -8065 Favorable =(1400*14.2)-(8100*13.8*15/60)